If Sammy can mow a lawn in 3 hours and Laura can mow the same lawn in 4 hours, how long will it take them if they work together?
Answer by richwmiller(17219)

x/3+x/4=1
4x+3x=12
7x=12
x=12/7 =1 5/7 hrs
about 1 hr 43 minutes